OpenStudy (anonymous):

What is the length of the hypotenuse of a right triangle with leg lengths of 8 inches and 15 inches?

OpenStudy (anonymous):

hypotenuse^2 = 8^2 + 15^2 hypotenuse^2 = 64 + 225 = 289 = 17^2 therefore hypotenuse = 17 inches

OpenStudy (anonymous):

use the pythagorean theorem c square= a square = b square c square= 8 square + 15 square c square= 64+225 c square= 289 c= square root of 289 c=17 ( c stands for hypotenuse a&b stands for length)